### v1

# D2USt3R: Enhancing 3D Reconstruction with 4D Pointmaps for Dynamic Scenes

Abstract
We address the task of 3D reconstruction in dynamic scenes, where object motions degrade the quality of previous 3D pointmap regression methods, such as DUSt3R, originally designed for static 3D scene reconstruction. Although these methods provide an elegant and powerful solution in static settings, they struggle in the presence of dynamic motions that disrupt alignment based solely on camera poses. To overcome this, we propose D2USt3R that regresses 4D pointmaps that simultaneiously capture both static and dynamic 3D scene geometry in a feed-forward manner. By explicitly incorporating both spatial and temporal aspects, our approach successfully encapsulates spatio-temporal dense correspondence to the proposed 4D pointmaps, enhancing downstream tasks. Extensive experimental evaluations demonstrate that our proposed approach consistently achieves superior reconstruction performance across various datasets featuring complex motions.
Figure 1: Teaser. Given a pair of input views, our D2USt3R  accurately establishes dense correspondence not only in static regions but also in dynamic regions, enabling full reconstruction of a dynamic scene via our proposed 4D pointmap. As highlighted by red dots $\bullet$ in the pointmap, DUSt3R [42] and MonST3R [48] align pointmaps solely based on camera motion, causing corresponding 2D pixels to become misaligned in 3D space. We compare the cross-attention maps, established correspondence fields, and estimated depth maps produced by our D2USt3R against baseline methods.
Co-first authors.Co-corresponding authors.
1 Introduction
Recovering 3D scene geometry from images remains a central challenge in computer vision. Traditional approaches, such as Structure-from-Motion (SfM) [32] and Multi-View Stereo (MVS) [34], have achieved impressive results in this context. While these methods are originally designed for recovering precise 3D scene geometry, they often struggle with scenes that include dynamic objects, symmetries, imagery with minimal overlapping and textureless regions [25, 33, 3, 9].
Recent approaches [42, 21, 40] leverage deep learning to streamline the 3D reconstruction pipeline and enhance robustness. DUSt3R [42], as a pioneering method, introduced a unified learning-based framework for dense stereo 3D reconstruction by jointly addressing image matching, essential matrix estimation, and triangulation. Specifically, DUSt3R directly regresses 3D pointmaps that encode scene geometry, pixel-to-scene correspondences, and inter-view relationships, mitigating error accumulation typical in multi-stage pipelines. Despite its strengths in static scenarios, DUSt3R significantly struggles with dynamic scenes due to its rigidity assumption, as exemplified in Figure 1.
Dynamic scenes, prevalent in real-world scenarios, pose significant challenges in 3D scene reconstruction task, as object motion disrupts the camera pose-based alignment [42], causing misaligned correspondences and inaccurate depth estimates. Recent methods such as MonST3R [48] tried to address dynamic 3D scene reconstruction by training on dynamic scene video data collections, but still relied on rigidly warped pointmaps within a common coordinate frame. Consequently, these approaches suffer from compromised correspondence learning for dynamic objects, impairing depth accuracy and robust geometry recovery.
In this paper, we propose Dynamic Dense Stereo 3D Reconstruction (D2USt3R), a novel feed-forward framework that directly regresses 4D pointmaps, simultaneously accounting for both spatial structure and motion to enable more reliable 3D reconstruction of both static and dynamic scene regions. This is achieved by directly introducing and aiming to establish dense correspondence fields between dynamic objects into the training process, ensuring consistent alignment of pointmaps across both static and dynamic regions. Compared to MonST3R [48], our model effectively captures dense correspondences between frames and enhances depth accuracy for dynamic elements by treating correspondence and reconstruction in dynamic scenes as entangled components. Our approach maintains the original DUSt3R [42] architecture, simply introducing dynamic motion awareness through a novel training signal. Experimental results and visual comparisons clearly demonstrate the significant improvement in reconstruction accuracy provided by our method.
Our key contributions are summarized as follows:
  • We introduce a novel 3D reconstruction framework that explicitly models dynamic motion through dynamically aligned 4D pointmaps, enabling unified reconstruction of complex scenes.
  • To compensate for the missing direct 3D correspondences between dynamic objects, we propose a 3D alignment loss that effectively handles occlusions and motions.
  • D2USt3R achieves state-of-the-art performance across several downstream tasks, including multi-frame depth estimation as well as camera pose estimation, demonstrating superior results in dense 3D reconstruction of dynamic scenes.
Reconstruction Correspondence
Methods Static Dynamic Static Dynamic
DUSt3R [42]
MASt3R [21] $\triangle$
MonST3R [48]
D2USt3R(Ours)
Table 1: Comparison of our method with existing approaches. Our proposed method uniquely handles both reconstruction and matching tasks consistently across static and dynamic scenarios. Note that MASt3R [21] employs a matching head to match points; however, it operates under the assumption of a static scene.
2 Related Work
Figure 2: Cross-attention visualization of DUSt3R [42] on static and dynamic scenes. We present visualizations of the source image’s attention maps corresponding to the query point (highlighted in red on the target image), both at each individual layer and as an average across layers. Although DUSt3R [42] effectively captures geometric correspondences during 3D reconstruction, it struggles to establish correspondences in regions exhibiting dynamic motion. This issue stems from its training regime, which assumes that both frames are static and can be modeled solely with rigid camera motion, thereby neglecting the complexities introduced by dynamic scenes.
Figure 3: Comparison of cross-attention on dynamic scene. Although MonST3R [48] is trained on dynamic video, its training signal remains identical to that of DUSt3R [42]. As a result, it fails to achieve reliable alignment between frames, ultimately limiting its 3D reconstruction performance. In contrast, our D2USt3R successfully establishes correspondences between dynamic frames, and thus has a stronger ability to estimate 3D shapes from dynamic motions.
Per-scene 3D reconstruction.
Classical 3D reconstruction methods typically follow a multi-stage pipeline to recover scene geometry and camera parameters from a set of uncalibrated images. Prominent examples include Structure-from-Motion (SfM) [49], and Simultaneous Localization and Mapping (SLAM) [8]. SfM incrementally reconstructs sparse 3D points through feature matching and bundle adjustment where SLAM simultaneously estimates camera trajectories and builds sparse or semi-dense maps in real-time. Building upon SfM, ParticleSfM [50] incorporate particle-based trajectory modeling to track object motions, paving the way for reconstruction in dynamic scenes. Despite all of these approaches showing remarkable performance, they typically require long processing time and resources for scene-specific optimization and often struggle with error accumulations from multi-stage pipelines.
Learning-based static scene reconstruction.
Building on top of estabilished correspondence fields [4, 5, 14, 16, 15], various learning-based approaches [40, 24, 12, 36, 18] have been proposed to reconstruct static 3D scenes by learning strong 3D priors, representing the scene as point clouds [24, 13], meshes [12, 41], voxels [36, 6] and 3DGS [17]. Recently, DUSt3R [42] notably provides a unified, feed-forward pipeline for dense stereo matching, geometry estimation, and triangulation by directly regressing structured 3D pointmaps. Although DUSt3R significantly improves reconstruction quality and efficiency by reducing cumulative errors, DUSt3R is inherently designed for static scenes, limiting its effectiveness in scenarios involving dynamic components.
Learning-based dynamic scene reconstruction.
Dynamic scene reconstruction introduces additional complexities due to non-rigid transformations occurring across frames. Similarly to static scene reconstruction, several recent approaches [23, 48, 27] have employed learning-based methods to tackle dynamic scene reconstruction. Among these, MonST3R [48] directly fine-tunes DUSt3R using dynamic videos, enabling a feed-forward approach to reconstruct dynamic scenes. Despite demonstrating strong overall performance, MonST3R retains DUSt3R’s per-frame training paradigm and lacks explicit mechanisms for linking corresponding points across frames in dynamic scenes. As demonstrated in Table 1, while DUSt3R is trained specifically to align corresponding 3D points, MonST3R does not effectively preserve this alignment during finetuning for dynamic content. Consequently, dynamic regions can suffer from inconsistent depth estimations, arising from inadequate temporal constraints that fail to capture the intricate motion patterns of objects. Our method addresses this limitation by augmenting the existing feed-forward framework with motion-aware training objectives, explicitly enforcing consistent 3D point correspondences over time.
3 Preliminary - DUSt3R
Given a pair of input images $I^{1},I^{2}\in\mathbb{R}^{W\times H\times 3}$, DUSt3R [42] predicts a pair of 3D pointmaps $X^{1,1},X^{2,1}\in\mathbb{R}^{W\times H\times 3}$ for both images, each expressed in the camera coordinate system of $I^{1}$. To train the network in a supervised manner, ground-truth pointmaps for each image are defined in the coordinate space of the first camera. Specifically, given the camera intrinsics matrix $K\in\mathbb{R}^{3\times 3}$, world-to-camera pose matrices $P_{n},P_{m}\in\mathbb{R}^{4\times 4}$ for images $n$ and $m$, and a ground-truth depth map $D\in\mathbb{R}^{W\times H}$, the ground-truth pointmap is computed as $X^{n,m}=P_{m}P_{n}^{-1}h(K^{-1}D)$, where $h:(x,y,z)\mapsto(x,y,z,1)$ represents the transformation to homogeneous coordinates.
Using 3D pointmaps, DUSt3R learns its parameters by minimizing the Euclidean distance between the ground-truth pointmaps $\bar{X}^{1,1}$, $\bar{X}^{2,1}$ and the predicted pointmaps $X^{1,1}$, $X^{2,1}$ for two corresponding sets of valid pixels $\mathcal{D}^{1},\mathcal{D}^{2}\subseteq\{1\dotsc W\}\times\{1\dotsc H\}$ on which ground-truth defined using a regression loss defined as:
$\mathcal{L}_{\text{regr}}(v,i)=\left\|\frac{1}{z}X_{i}^{v,1}-\frac{1}{\bar{z}}\bar{X_{i}}^{v,1}\right\|,$ (1)
where $v\in\{1,2\}$ denotes the input views and $i\in\mathcal{D}^{v}$ denotes valid pixel positions. The scaling factors $z=\text{norm}(X^{1,1},X^{2,1})$ and $\bar{z}=\text{norm}(\bar{X}^{1,1},\bar{X}^{2,1})$ are computed using the normalization function:
$\text{norm}(X^{1},X^{2})=\frac{1}{|\mathcal{D}^{1}|+|\mathcal{D}^{2}|}\sum_{v\in\{1,2\}}\sum_{i\in\mathcal{D}^{v}}\|X^{v}_{i}\|.$ (2)
Additionally, DUSt3R incorporates a confidence score to learn to reject errorneously defined GT pointmaps, and it is included in the final loss function such that:
$\mathcal{L}_{\text{conf}}=\sum_{v\in\{1,2\}}\sum_{i\in\mathcal{D}^{v}}C_{i}^{v,1}\mathcal{L}_{\text{regr}}(v,i)-\alpha\log C_{i}^{v,1}.$ (3)
4 Methodology
Figure 4: Construction of alignment loss defined at static and dynamic regions. We propose a pipeline for generating a 4D pointmap while explicitly addressing occlusions caused by dynamic regions. (a) From the input images, we obtain optical flow refined via cycle consistency checks and derive a dynamic mask $M_{\text{dyn}}$. To align image $I^{2}$ with image $I^{1}$, we utilize: 1) the camera pose to align static regions in $I^{2}$, and 2) optical flow to align dynamic regions. The alignment process is conducted specifically within regions corresponding to the colored pixels. Through this process, we construct a 4D pointmap, enabling alignment in 3D space for all corresponding 2D pixels.
4.1 Motivation and overview
As visualized in Figure 1, DUSt3R [42] learns to predict 3D pointmaps primarily defined in static regions, enabling accurate and robust 3D reconstruction through precise stereo correspondences. Motivated by ZeroCo’s [1] finding that the cross-attention maps of DUSt3R inherently encode geometric correspondence information, we visualize the cross-attention maps for both static and dynamic scenes in Figure 2. However, the underlying assumption that reliable pointmaps exist only in static regions limits its applicability to dynamic scenes. To compensate for, recent methods, such as MonST3R [48], have attempted to address dynamic scenes by extending DUSt3R’s framework to support 4D reconstruction through augmented training datasets that include dynamic scenarios. Nevertheless, this approach overlooks the explicit correspondence between dynamic objects, resulting in inaccurate depth estimations. This omission impairs the ability to leverage dynamic objects as anchor points, which could otherwise reinforce the spatial structure and improve depth estimation of nearby static regions, ultimately hindering overall 3D reconstruction performance. Figure 3 visualizes these issues through attention maps, highlighting noisy correspondences that degrade reconstruction accuracy.
Motivated by these limitations, we propose extending 3D pointmaps to 4D pointmaps to explicitly encode motion dynamics. Our method ensures accurate alignment not only in static regions but also among dynamic objects. In the following sections, we describe our proposed 4D pointmap representation and outline our training procedure.
4.2 Occlusion and dynamic masks preparation
We find that existing datasets often do not include 3D correspondences between dynamic objects, which poses additional challenges in learning them. To alleviate this, we first employ off-the-shelf optical flow [44] to find dense 2D correspondences across image pairs. Note that if the ground-truth optical flows are available in the dataset, we instead use them. For the scenarios where occlusions are more frequent due to large camera baselines, we perform forward-backward consistency check [38, 47] to additionally obtain occlusion masks. This process is defined as follows:
$p^{\prime}_{2}=p_{1}+\mathbf{f}(p_{1}),\quad p^{\prime}_{1}=p^{\prime}_{2}+\mathbf{b}(p^{\prime}_{2}),\quad$ (4)
${M}_{\mathrm{occ}}=|p^{\prime}_{1}-p_{1}|>t,$
where $p_{i}$ denotes a pixel in image $I^{i}$, $\mathbf{f}$ and $\mathbf{b}$ represent forward and backward optical flows, respectively, and $t$ is an occlusion threshold.
While we empirically find that occlusion mask provides reliable dense correspondences for supervision, due to frequent large changes in camera and dynamic objects motions, we find it helpful to introduce a dynamic mask ${M}_{\mathrm{dyn}}$, to encourage stable learning process by informing the network the regions of dynamic objects within the images. Specifically, the dynamic mask ${M}_{\mathrm{dyn}}$ is computed by comparing optical flow $\mathbf{f}$ with the expected flow induced purely by camera motion $\mathbf{f}_{\text{cam}}$. Given the depth map $D$, intrinsics matrix $K$, relative rotation and translation $R,T$, and pixel coordinates $p$, we define:
$\mathbf{f}_{\text{cam}}$ $=\pi(DKRK^{-1}p+KT)-p,$ (5)
$\mathcal{\mathcal{missing}}{M}_{\mathrm{dyn}}$ $=[\|\mathbf{f}_{\text{cam}}-\mathbf{f}\|<\tau],$
where $\pi:(x,y,z)\mapsto(x/z,y/z)$ denotes the projection operation, and $\tau$ serves as the dynamic threshold.
Dataset Domain ##\## of frames ##\## of Scenes Dynamics Dynamic mask Optical flow Ratio
Blinkvision Outdoor [22] Outdoors 6k 23 Realistic 38.75%
Blinkvision Indoor [22] Indoors 6k 24 Realistic 23.75%
PointOdyssey [51] Indoors & Outdoors 200k 131 Realistic 12.5%
TartanAir [43] Indoors & Outdoors 1000k 163 None 12.5%
Spring [28] Outdoors 6k 37 Realistic 12.5%
Table 2: Training datasets utilized for fine-tuning the pretrained DUSt3R [42] on dynamic scenes. All datasets consists of synthetic scenes and provide both camera pose and depth information, with most containing dynamic objects. For the PointOdyssey dataset, we excluded scenes containing smoke or those with inaccurate segmentation masks. For the Blinkvision Outdoor dataset, scenes with incomplete dataset annotations were omitted. Details regarding each scene are documented in the supplementary material A. Training dataset.
4.3 4D pointmaps regression
Pointmap alignment in static regions.
The regression loss in DUSt3R inherently aligns 3D pointmaps using camera pose alone. To ensure alignment focuses solely on static regions within image $I^{2}$, we employ the dynamic mask ${M}_{\mathrm{dyn}}$ and restrict the computation of the regression loss accordingly. Thus, we modify the regression loss $\mathcal{L}_{\text{regr}}$ as follows:
$\mathcal{L}_{\text{regr}}(1,i)$ $=\left\|\frac{1}{z}X_{i}^{1,1}-\frac{1}{\bar{z}}\bar{X_{i}}^{1,1}\right\|,$ (6)
$\mathcal{L}_{\text{regr}}(2,i)$ $=\left(1-M^{2}_{\text{dyn},i}\right)\left\|\frac{1}{z}X_{i}^{2,1}-\frac{1}{\bar{z}}\bar{X_{i}}^{2,1}\right\|.$
To account for errorneously defined GT pointmaps, we additionally introduce a confidence-aware loss in static regions, $\mathcal{L}_{\text{static}}$, to incorporate uncertainties into the alignment process:
$\mathcal{L}_{\text{static}}=\sum_{v\in\{1,2\}}\sum_{i\in\mathcal{D}^{v}}C_{i}^{v,1}\mathcal{L}_{\text{regr}}(v,i)-\alpha\log C_{i}^{v,1}.$ (7)
Pointmap alignment in dynamic regions.
To align dynamic objects from $I^{2}$ to $I^{1}$, we introduce a dynamic alignment loss that leverages both the occlusion mask ${M}_{\mathrm{occ}}$ and the dynamic mask ${M}_{\mathrm{dyn}}$ to effectively address occlusions and motion. As illustrated in Figure 4, these masks are computed in a dedicated pipeline. This method ensures that when points from the second view $I^{2}$, are transformed into the coordinate system of $I^{1}$, they accurately correspond to the temporal state of the first view. In line with our regression loss, we further incorporate confidence estimates to define a confidence-aware alignment loss. The dynamic alignment loss is formulated as follows:
$\mathcal{L}_{\text{dyn}}=$ (8)
$\sum_{i\in\mathcal{D}^{2}}M^{2}_{\mathrm{dyn},i}(1-M^{2}_{\mathrm{occ},i})C_{i}^{2,1}\left\|\frac{1}{\bar{z}_{1}}\bar{X}_{i+\mathbf{b}(i)}^{1,1}-\frac{1}{z_{1}}X_{i}^{2,1}\right\|$
$+\sum_{i\in\mathcal{D}^{1}}M^{1}_{\mathrm{dyn},i}(1-M^{1}_{\mathrm{occ},i})C_{i}^{1,2}\left\|\frac{1}{\bar{z}_{2}}\bar{X}_{i+\mathbf{f}(i)}^{2,2}-\frac{1}{z_{2}}X_{i}^{1,2}\right\|.$
We leverage optical flow $\mathbf{f}$ to establish dense correspondences between $I^{1}$ and $I^{2}$, and its reverse direction $\mathbf{b}$. This mechanism ensures that both views are aligned within a unified coordinate framework. The first term in our loss function enforces the alignment of the pointmap in the coordinate space of the camera system associated with $I^{1}$, while the second term introduces a symmetric constraint by aligning the points when the roles of the views are swapped.
Figure 5: Qualitative results for pointmap reconstruction. We qualitatively compare the 3D pointmap of D2USt3R against other pointmap regression models. All visualizations presents per-pixel pointmaps without applying confidence thresholding. It is notable that both DUSt3R [42] and MonST3R [48] struggle to accurately reconstruct scenes that include dynamic elements. We find that inaccurately established correspondence fields between dynamic regions negatively affect the overall reconstruction performance.
Figure 6: Depth estimation qualitative results.
Figure 7: Visualization of correspondences given a pair of images. We show that our method can find accurate correspondences between dynamic objects.
Final Objective.
Our final objective function is defined as following:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{static}}+\mathcal{L}_{\text{dyn}}$ (9)
This combined loss function enables our model to capture precise 3D geometry and robust correspondences in dynamic scenes while retaining DUSt3R’s proven benefits in static regions.
4.4 Additional heads for downstream task
Dynamic mask head.
Since our model implicitly understands the mask’s location, we further regress a dynamic mask using an additional head. We predict a single-channel logit, $\hat{M}_{\text{dyn}}$, using a DPT head [31] in a manner similar to the pointmap regression head. To train this head, we employ the following cross-entropy loss:
$\mathcal{L}_{\text{mask}}=-\frac{1}{|\mathcal{D}_{\text{all}}|}$ $\sum_{i\in\mathcal{D}_{\text{all}}}\left[M_{\text{dyn},i}\log(\sigma(\hat{M}_{\text{dyn},i}))\right.$ (10)
$+\left.(1-M_{\text{dyn},i})\log(1-\sigma(\hat{M}_{\text{dyn},i}))\right],$
where $\sigma(\cdot)$ denotes the sigmoid function and $\mathcal{D}_{\text{all}}$ denotes the set of all pixels.
Optical flow head.
For accurate optical flow estimation, we incorporate an additional flow head built on the RAFT architecture [39]. Inspired by insights from ZeroCo [1], our flow head utilizes cross-attention maps rather than conventional 4D correlation volumes. Empirically, we find that this helps to further enhances the accuracy of estiamted flow maps. We train this head using the Mixture-of-Laplace loss [44] as the supervision signal.
5 Experiments
Category Methods TUM-Dynamics Bonn Sintel KITTI
All Dynamic All Dynamic All Dynamic All
[…]
Single-frame depth DepthAnythingv2 [46] 0.098 89.0 - - 0.073 93.8 - - 0.336 55.6 - - 0.069 93.7
Marigold [19] 0.205 72.3 - - 0.066 96.4 - - 0.623 50.5 - - 0.104 89.9
Multi-frame depth DUSt3R [42] 0.176 76.5 0.221 71.3 0.135 82.4 0.127 83.7 0.370 58.5 0.672 54.9 0.076 93.6
MASt3R [21] 0.165 79.0 0.199 73.8 0.183 77.5 0.167 79.5 0.330 57.3 0.528 54.4 0.050 96.8
MonST3R [48] 0.145 81.2 0.152 79.2 0.068 94.4 0.066 94.9 0.345 56.2 0.525 46.9 0.070 95.0
MonST3R* [48] 0.159 81.0 0.181 76.5 0.076 93.9 0.071 94.4 0.349 52.5 0.565 36.9 0.103 90.9
D2USt3R(Ours) 0.142 83.9 0.148 82.9 0.060 95.8 0.059 95.7 0.324 57.5 0.568 48.0 0.104 90.7
Table 3: Multi-frame depth estimation results. We compare multi-frame depth for both the entire scene and dynamic regions separately. The comparison for dynamic regions is conducted only when the dynamic parts are identifiable. *: Reproduced with same dataset as Ours.
Methods Bonn Sintel KITTI NYU-v2 TUM-Dynamics
[…]
DUSt3R [42] 0.141 82.5 0.424 58.7 0.112 86.3 0.080 90.7 0.176 76.8
MASt3R [21] 0.142 82.0 0.354 57.9 0.076 93.2 0.129 84.9 0.160 78.7
MonST3R [48] 0.076 93.9 0.345 56.5 0.101 89.3 0.091 88.8 0.147 81.1
MonST3R* [48] 0.083 93.6 0.387 50.6 0.143 85.0 0.084 90.1 0.163 79.1
D2USt3R(Ours) 0.065 95.2 0.340 58.4 0.131 86.2 0.085 90.1 0.150 82.9
Table 4: Single-frame depth estimation results. *: Reproduced with same dataset as Ours.
Methods Sintel TUM-Dynamics ScanNet
Rotation Translation Rotation Translation Rotation Translation
[…]
DUSt3R [42] 6.15 4.51 0.29 0.26 2.36 0.98 0.013 0.01 0.74 0.54 0.11 0.08
MASt3R [21] 4.71 3.40 0.23 0.19 2.83 1.13 0.06 0.03 0.85 0.64 0.05 0.04
MonST3R [48] 4.90 2.30 0.26 0.22 1.88 1.39 0.019 0.01 0.94 0.79 0.10 0.08
MonST3R* [48] 8.50 2.61 0.27 0.23 1.76 1.40 0.02 0.01 0.74 0.58 0.10 0.08
D2USt3R(Ours) 6.96 2.67 0.26 0.22 1.80 1.41 0.03 0.02 0.75 0.57 0.08 0.06
Table 5: Camera pose estimation results. *: Reproduced with same dataset as Ours.
Methods Sintel-Clean Sintel-Final KITTI
DUSt3R [42] 30.96 35.11 14.19
MASt3R [21] 39.37 39.50 13.27
MonST3R [48] 38.47 41.92 14.91
MonST3R* [48] 37.47 40.58 14.58
D2USt3R(Ours) 16.19 25.31 8.91
Croco-Flow [45] 3.31 4.28 13.24
SEA-RAFT [44] 5.21 13.18 4.43
D2USt3R + Flow head 9.25 12.77 3.57
Table 6: Evaluation of pointmap alignment accuracy in dynamic objects. We evaluate the Flow End-Point Error (EPE) $\downarrow$ on the Sintel and KITTI datasets. Note that Croco-Flow [45] was grayed out in the Sintel evaluation since it was trained on the Sintel dataset. *: Reproduced with same dataset as Ours.
5.1 Experimental setup
Implementation details.
Building on top of DUSt3R [42], we freeze the encoder and fine-tune only the decoder and the DPT head [31], as done similarly by MonST3R. For each epoch, we randomly sample 20,000 image pairs and the network is trained for 50 epochs. We use the AdamW optimizer [26] with an initial learning rate of 5e-5. We train with 4 NVIDIA RTX 6000 GPUs, with a batch size of 4 images per GPU and gradient accumulation steps set to 2. More details can be found in .
Training datasets.
As shown in Table 2, we train D2USt3R  on multiple datasets, including BlinkVision Outdoor [22], BlinkVision Indoor [22], Spring [28], PointOdyssey [51], and TartanAir [43]. Each epoch consisted of sampling 7,750, 4,750, 2,500, 2,500, and 2,500 pairs, respectively. Additionally, we perform random sampling with temporal strides varying from 1 to 9 to account for large camera motions and dynamic scenarios.
Baselines.
Following [42], we evaluate our method on depth estimation and camera pose estimation. We additionally evaluate pointmap alignment accuracy in dynamic regions. We compare our method against existing state-of-the-art pointmap regression models, specifically DUSt3R [42], MASt3R [21], and MonST3R [48]. Furthermore, to ensure fair comparisons and demonstrate the effectiveness of our approach for predicting the 4D pointmap, we trained a variant of MonST3R, termed MonST3R*, under the same setup as ours.
Evaluation setup.
For multi-frame depth estimation, we evaluate on the TUM-Dynamics [37], Bonn [29], Sintel [2], KITTI [11], and ScanNet [7] datasets using image pairs with source frames offset from the target by strides of 1, 3, 5, 7, and 9 frames. We assess performance over the entire scene and exclusively on dynamic regions when dynamic masks are available. For single-frame depth estimation, we evaluate on the Bonn [29], Sintel [2], KITTI [10], NYU-v2 [35], and TUM-Dynamics [37] datasets. In both settings, we follow the affine-invariant depth evaluation protocol, reporting the Absolute Relative Error (AbsRel) and the percentage of inlier points ($\delta_{1}$, where $\delta<1.25$).
5.2 Experimental results
Depth estimation.
In this experiment, we evaluate our method and compare with existing methods on single/multi-frame depth estimation. The results are summarized in Table 3 and  4. We also show qualitative comparisons in Figure 5 and Figure 6. From the results, we observe that D2USt3R outperforms other methods overall. However, our performance on KITTI is somewhat limited, likely due to the absence of driving scenes in our training data, which may have disadvantaged our model. Nonetheless, we compare with MonST3R, which was trained with the same training datasets as ours, and find that the performance is comparable. Finally, we highlight in Figure 5 that our approach consistently predicts accurate depth for dynamic human subjects.
Camera pose estimation.
In Table 5, we evaluate camera pose performance of our model on the Sintel [2], TUM-dynamics [37] and ScanNet [7] datasets. Note that we use the 2D–3D matching between the prediction $X^{2,1}$ and the original coordinates to obtain camera poses via PNP-RANSAC.
Dynamic alignment.
In Table 6, we evaluate pointmap alignment accuracy in dynamic objects. For this, we use Sintel [2] and KITTI [10]. Since we directly obtain aligned pointmaps even in dynamic objects, we can easily derive 2D-2D matching points by computing the difference between $X^{2,1}$ and $X^{1,1}$. From the results, we find that our method outperforms other baselines, which is further supported in Figure 7, where accurately captured correspondences between objects in different time step frames are observed. This gap is further broadened when we leverage an additional flow head, as shown in Table 6, where ours with the flow head outperforms the state-of-the-art SEA-RAFT [44]. We provide qualitative examples in Figure 8.
Dynamic mask.
In this experiment, we show that using dynamic mask head, our method reliably predicts dynamic regions. We show visualizations in Figure 9, where the dynamic mask head effectively segments dynamic objects across diverse in-the-wild scenarios.
Methods TUM-Dynamics Bonn
[…]
Full finetune 0.161 77.0 0.081 91.9
Finetune decoder & head 0.142 83.9 0.060 95.8
Table 7: Ablation on training strategy.
5.3 Ablation study
While we have already demonstrated the effectiveness of our proposed 3D alignment loss and 4D pointmaps are validated in previous experiments, we additionally explore the impacts of different finetuning strategies. The results are shown in Table 7. From the results, compared to full fine-tuning, partial fine-tuning of only the decoder and the downstream head provides more competitive results. We thus only train the decoder and head in this work.
Figure 8: Visualization of predicted optical flow.
Figure 9: Visualization of predicted dynamic mask.
6 Conclusion
In this paper, we have introduced a novel approach for 3D dynamic scene reconstruction, featuring a simple yet effective extension to existing pointmap representations to accommodate dynamic motions. Our proposed method significantly enhances the quality of 3D reconstruction in dynamic environments. We evaluated our approach comprehensively across tasks including depth estimation, camera pose estimation, and 3D point alignment. Experimental results demonstrate that our method outperforms existing approaches on real-world, large-scale datasets, achieving new state-of-the-art performance.
Supplementary Material
A. Training dataset
Dataset Excluded Scenes
PointOdyssey animal_s, animal_smoke_, animal1_s_, animal1_s, animal2_s, animal3_s, animal4_s, animal6_s,
cab_e_ego2, cab_e1_3rd, cab_e1_ego2, cab_h_bench_3rd, cab_h_bench_ego2, character0_f, character0_f2,
character3_f, character4_, character5_, character6, cnb_dlab_0215_3rd, cnb_dlab_0215_ego1,
cnb_dlab_0225_3rd, cnb_dlab_0225_ego1, dancingroom_3rd, human_in_scene, kg, r5_new_f,
scene_d78_0318_3rd, scene_d78_0318_ego1, scene_d78_0318_ego2, scene_j716_3rd, scene_j716_ego1,
scene_j716_ego2, scene1_0129, seminar_h52_ego1
Blinkvision Outdoor outdoor_train_autopilot_tree_01, outdoor_train_autopilot_tree_02,
outdoor_train_autopilot_tree_03, outdoor_train_track_animal_people
Table 8: List of excluded scenes from the datasets.
We excluded invalid scenes from the PointOdyssey, Blinkvision Outdoor utilized in Table 2. Specifically, in the PointOdyssey dataset, we removed scenes containing smoke or inaccurate annotations. Additionally, in the Blinkvision Outdoor dataset, we excluded scenes with incomplete dataset annotations. The excluded scenes are listed in Table 8.
B. Bullet time reconstruction for dynamic video input
Figure 10: Visualization of bullet-time reconstruction from a long sequence of dynamic video inputs. We visualize the reconstruction of bullet-time view (20 frames) from a dynamic video input consisting of 40 frames. Since MonST3R is incapable of dynamic alignment, it predicts depth independently at each timestep, similar to monocular depth estimation, resulting in the reconstruction as a sequence of 3D pointmaps.
Owing to dynamic alignment and 4D pointmaps, we can aggregate and render a highly dynamic video input into a single, coherent bullet-time view. As illustrated in Figure 10, a video consisting of large input frames can be aligned into an intermediate bullet-time representation. This approach enables static reconstruction even when the input video includes dynamic motion. Consequently, it becomes feasible to directly train a 3D Gaussian splatting [20] on scenes containing moving people in landmarks or dynamics that are challenging to render using conventional methods.
C. Additional results
Figure 11: Additional qualitative results for pointmap reconstruction. We qualitatively compare the 3D pointmap of D2USt3R against MonST3R. All visualizations presents per-pixel pointmaps without applying confidence thresholding.
Figure 12: Additional qualitative results for depth estimation on Sintel dataset.
Figure 13: Qualitative results for optical flow estimation on DAVIS dataset.
Figure 14: Qualitative results for optical flow estimation on KITTI dataset.
In Figure 11 and Figure 12, we additionally present qualitative results for pointmap reconstruction and depth estimation. Utilizing our 4D pointmaps, dynamic elements are well-aligned, thereby enhancing both depth estimation and 3D reconstruction quality. In Figure 13 and Figure 14, we present qualitative results for optical flow on DAVIS [30] and KITTI [11] datasets. We observed that D2USt3R is capable of accurately predicting optical flow without relying on any dedicated optical flow module.